OpenStudy (anonymous):

consider a parabola, P, that is congruent to y= x^2 and with vertex (2,-4). find the equation of a new parabola that results if P is translated 3 units to the right and translated 1 unit down.

OpenStudy (anonymous):

y+5 = 0.25(x-5)^2

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

see the eqn of the parabola is (y-k)=c(x-h)^2

OpenStudy (anonymous):

translate the vertex 3 to right and one down it b ecomes (h,K)

OpenStudy (anonymous):

since it is congruent to y=x^2, c is the same for both ie..1/4